Two balls are drawn in succession without replacement from an urn containing 6 green balls and 6 yellow balls. Let Y be the random variable representing the number of yellow balls. Find the values of the random variable Y

1
Expert's answer
2021-12-20T07:10:15-0500

Given,

6 yellow balls and 6 green balls.

Total=12


Z is a random variable So value of Z can be 0,1 or 2 as two balls are drawn in succession.


When there is 0 yellow balls So both balls are green.


"P(Z=0)=\dfrac{^6C_2}{^{12}C_2}=0.23"



Now If there is one yellow ball ,This can be done in two ways either green first then yellow or vice versa.


"P(Z=1)=\dfrac{6}{12}\times \dfrac{6}{11}+\dfrac{6}{12}\times \dfrac{6}{11}=0.545"


When both balls are yellow.


"P(Z=2)=\dfrac{^6C_2}{^{12}C_2}=0.23"


Probability Distribution table is-